OpenStudy (anonymous):

Use the distance formula to find the distance between the two points . (-3,2) (1,-3) .

OpenStudy (anonymous):

\[\sqrt{(x2-x1)+(y2-y1)}\]\[\sqrt{(1+3)+(2+3)}\]\[\sqrt{4+5}\]\[\sqrt{9}\] =3